OpenStudy (anonymous):

what is the domain of sqrt (6x - 6) and sqrt (-5x -6)

OpenStudy (danmac0710):

You can only find SQRTs of numbers greater than or equal to zero so therefore your (6x-6) and (-5x-6) have to be obey this. SOlve those simple inequalities and you have your domain. Good luck.

OpenStudy (danmac0710):

Do you know how to solve linear inequalities?

OpenStudy (anonymous):

unfortunately not

OpenStudy (danmac0710):

Right: [6x-6\ge0\] \[6x \ge6\] \[x \ge1\] That's your domain...

OpenStudy (danmac0710):

Hang on that looks weird.... first line should be: \[6x-6\ge0\]

OpenStudy (anonymous):

thanks how do i write that in interval notation

OpenStudy (danmac0710):

\[x \in[1,\infty[\]

OpenStudy (danmac0710):

Actually you can just use: \[x \ge1\] That is also correct interval notation.

OpenStudy (anonymous):

ok about the other one

OpenStudy (anonymous):

do i set it up as -5x-6= 0 the i add 6

OpenStudy (anonymous):

and it now looks like this -5x=6

OpenStudy (danmac0710):

If you are needing to state domains of functions you probably need to learn how to solve linear inequalities as well. Otherwise your exam will go pear-shaped!

OpenStudy (anonymous):

x=6/-5

OpenStudy (danmac0710):

That is right

OpenStudy (danmac0710):

except you need to use the greater than or equal to sign rather than just equals

OpenStudy (anonymous):

my software isnt accpeting the second one

OpenStudy (danmac0710):

Oh.... try writing is as -1.2

OpenStudy (danmac0710):

6 fifths but negative

OpenStudy (anonymous):

its still not

OpenStudy (anonymous):

did i do something wrong somewehere

OpenStudy (anonymous):

hey can you help me with this one\[1/\sqrt[4]{x ^{2}-2x}\]

OpenStudy (anonymous):

what is the domain